OpenStudy (anonymous):

integrating from sinx to 1 (t^2 f(t)dt=1-sinx for all xE R ,then f(1/sq.root(3))

OpenStudy (anonymous):

Try using the Equation button in the bottom.

OpenStudy (anonymous):

Is it \[\int\limits_{\sin x}^{1}t^2f(t) dt=1-\sin x?\]

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Just tell me if what I wrote there is what you're asking about :)

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):

Good. Do you know the fundamental theorem of calculus?

OpenStudy (anonymous):

i dunn mayb

OpenStudy (anonymous):

Hmm. I will start by taking the derivative of both sides. Taking the derivative of the integral can be done using: \[\int\limits_{a}^{g(x)}f(t)dt=f(g(x)).g'(x)\]

OpenStudy (anonymous):

Sorry write before the integral d/dx

OpenStudy (anonymous):

but y

OpenStudy (anonymous):

oh i gotta

OpenStudy (anonymous):

\[{d \over dx}\int\limits_{\sin x}^{1}t^2f(t)dt={d \over dx}(1-\sin x)\]

OpenStudy (anonymous):

You are leaving?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You got the formula I wrote above?

OpenStudy (anonymous):

yaa but then with this we hav to find f(\[ 1/\sqrt{3}\])

OpenStudy (anonymous):

I know.. We have to find f(x) first.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

It will take just two more steps.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\implies -(\sin^2 x)f(\sin x)\cos x=-\cos x \implies \cos x(- \sin^2 x f(\sin x)+1)=0\]

OpenStudy (anonymous):

It follows that: cos x=0 or \[\sin^2 x f(\sin x)=1 \implies f(\sin x)={1 \over \sin^2 x}\]

OpenStudy (anonymous):

Now to find the value of f(1/sqrt(3)).. put x=pi/3

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Sorry.. not pi/3, just substitute x=sin^-1(1/sqrt(3))you will get:

OpenStudy (anonymous):

bot acc to formula g(x) =1 then y we substitute sinx

OpenStudy (anonymous):

\[f(\sin(\sin^{-1}(1/\sqrt3))=f(1/\sqrt3)={1 \over \sin^2(\sin^{-1}{1\over \sqrt3})}={1 \over 1/3}=3\]

OpenStudy (anonymous):

Do you have the answer?

OpenStudy (anonymous):

you know that sin(arcsinx)=x, right? because arcsin just "undoes" the sin since it's its inverse.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Ok.. that's what I got and showed above. I hope the steps are clear, and make sense to you.

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

r u teacher

OpenStudy (anonymous):

No. I am a student.

OpenStudy (anonymous):

quit intelligent

OpenStudy (anonymous):

Thanks!! :)

OpenStudy (anonymous):

passed out 12 this yr

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

hav u passed out 12 this yr?

OpenStudy (anonymous):

hav u passed out 12 this yr?

OpenStudy (anonymous):

I don't know what that means.